OpenStudy (anonymous):

What is the value of the expression 8–2? The 2 is a small number.

OpenStudy (steve816):

8 - 2 = 6

OpenStudy (anonymous):

stop.

OpenStudy (steve816):

?

OpenStudy (anonymous):

Bump

OpenStudy (anonymous):

@marco26

OpenStudy (anonymous):

Can you help?

OpenStudy (anonymous):

Need help anyone?

OpenStudy (anonymous):

I need help yo

OpenStudy (anonymous):

@Black_ninja123

imqwerty (imqwerty):

ok so 8-2 means that ur subtracting 2 from 8 for example if u got 8 chocolates and i came and stole 2 of them then how many wuld be left with u :)

OpenStudy (anonymous):

No no no no.

OpenStudy (anonymous):

8-2 the 2 is a small number that is high

imqwerty (imqwerty):

nononono

imqwerty (imqwerty):

u jst want the answer :P

OpenStudy (anonymous):

Look What is the value of the expression 8–2?

imqwerty (imqwerty):

its 100

OpenStudy (anonymous):

The 2 is a small number

imqwerty (imqwerty):

yes 2 is smaller so what

OpenStudy (anonymous):

Look stay here I'll post a pic of what the 2 looks like

imqwerty (imqwerty):

ok :) please post a HDR pic cause i can't see what 2 looks like vry well :)

OpenStudy (anonymous):

http://www.picpaste.com/Capture-bM6Oz5lT.PNG

OpenStudy (anonymous):

See?

OpenStudy (anonymous):

Can you help?

imqwerty (imqwerty):

oh :) its\[8^\left( -2 \right)\] ok :) lol i thought u were kiddin XD

OpenStudy (drummergirl3):

lol

OpenStudy (anonymous):

How did you make the 2 look like taht

OpenStudy (anonymous):

So it will be alot easier for me to explain

imqwerty (imqwerty):

ok try to use this identity- \[a^\left( -b \right)=\frac{ 1 }{ a^b }\] i used the equation tool to do that :)

OpenStudy (anonymous):

What is it under?

OpenStudy (rushwr):

\[8^{-2} = \frac{ 1 }{ 8^{2} } = \frac{ 1 }{ 64}\]

OpenStudy (anonymous):

Rush

OpenStudy (anonymous):

Can you help me out with 4 more? WIll give you medals

OpenStudy (anonymous):

@Rushwr

imqwerty (imqwerty):

( ͡° ͜ʖ ͡°)

OpenStudy (anonymous):

Qwerty where is the smal letter thing in equations

OpenStudy (rushwr):

yeah yeah @dredayAJO

OpenStudy (anonymous):

Rushwr

OpenStudy (anonymous):

@Rushwr

OpenStudy (rushwr):

yh

OpenStudy (anonymous):

Can you help with 4 questions?

OpenStudy (anonymous):

http://picpaste.com/upload.php